OpenStudy (anonymous):

A soccer team played 160 games and won 65 percent of them. How many games did it win?

OpenStudy (anonymous):

65/100=x/160 x=104

OpenStudy (anonymous):

Why bring algebra into this?? You probably hate algebra anyway. To find a % of something just change the % to a decimal and multiply. 65% of 160 is just 0.65 x 160

OpenStudy (anonymous):

thanks! (: